OpenStudy (anonymous):

Solve for x: x + 3 = 4(3x - 1) A.7/11 B.5/11 C.-1/11 D.4/11 E.None of the above

OpenStudy (anonymous):

x+3=12x-4 3+4=11x x=7/11 so A